OpenStudy (anonymous):

i need help!! find the exact value of sin(2cos^-1(7/25))

OpenStudy (anonymous):

sin[2sin^-1(7/25) ] = sin A = 336/625

OpenStudy (anonymous):

How? lol. I would like to know the steps so i can learn it

OpenStudy (anonymous):

sin [ 2 sin^-1(7/25) ] let 2sin^-1(7/25) = A sin^-1(7/25) = A/2 sin(A/2) = 7/25 cos(A/2) = √(1 - 49/625) = √576/625 = 24/25 sin A = 2sin(A/2) cos(A/2) = 2(7/25)(24/25) = 336 /625 sin[2sin^-1(7/25) ] = sin A = 336/625